OpenStudy (anonymous):

A corner store sells two kinds of baked goods: cakes and pies. A cake costs $5 and a pie costs $7. In one day, the store sold 15 baked goods for a total of $91. How many cakes did they sell?

OpenStudy (anonymous):

use the equations: let x=cake y=pie (1) x+y=15 5x+7y=91

OpenStudy (anonymous):

how many cakes

OpenStudy (anonymous):

there are 8 cakes. .

OpenStudy (anonymous):

this answer was wrong